the car that will tow the trailer is restricted to pull a mass of not more than 500kg.The mass of each canned fish is 440g and the mass of each empty box is 27g. The school then decided to buy only 15 boxes of canned fish needed for a month. Determine the total mass in kg of each box containing 39 canned fish


1
Expert's answer
2021-09-06T16:21:46-0400

The mass of each canned fish is "0.440\ kg" and the mass of each empty box is "0.027\ kg."

Each box contains 39 canned fish.

The total mass in kg of each box


"m_1=39(0.440\ kg)+0.027\ kg=17.187\ kg"

The total mass in kg of 15 boxes


"m=15m_1=15(17.187\ kg)=257.805\ kg<500\ kg"